Please provide solution and explanation for the flwg:
X/180-x = 1/8
Don't understand how to eliminate fractions.
Thank you

Found 2 solutions by Earlsdon, Scorinitron:
Answer by Earlsdon(6294)

flwg???
Solve: x/(180 - x) = 1/8 or (x/180) - x = 1/8? It's probably x/(180-x) = 1/8
x/(180-x) = 1/8 Multiply both sides by (180-x)
x = (180-x)/8 Multiply both sides by 8.
8x = 180-x Add x to both sides.
9x = 180. Divide both sides by 9.
x = 20
Check:
20/(180-20) = 20/160 = 1/8

Answer by Scorinitron(19)

(x)/(180)-x=(1)/(8)

Find the LCD (least common denominator) of (x)/(180)-x+(1)/(8).

Least common denominator: 360
Multiply each term in the equation by 360 in order to remove all the
denominators from the equation.
(x)/(180)*360-x*360=(1)/(8)*360
Cancel the common factor of 180 in the denominator of the first term (x)/(180)

and the second term 360.

(x)/(180)*360^(2)-x*360=(1)/(8)*360

Reduce the expression by removing the common factor of 180 in the denominator
of the first term (x)/(180) and the second term 360.

x*2-x*360=(1)/(8)*360

Multiply x by 2 to get 2x.

2x-x*360=(1)/(8)*360


Multiply -x by 360 to get -360x.

2x-360x=(1)/(8)*360

According to the distributive property, for any numbers a, b, and c,

a(b+c)=ab+ac and (b+c)a=ba+ca. Here, x is a factor of both 2x and -360x.

(2-360)x=(1)/(8)*360

To add integers with different signs, subtract their absolute values and give the result the same sign as the integer with the greater absolute value. In this example, subtract the absolute values of 2 and -360 and give the result the same sign as the integer with the greater absolute value.

(-358)x=(1)/(8)*360


Remove the parentheses.

-358x=(1)/(8)*360

Cancel the common factor of 8 in the denominator of the first term (1)/(8) and the second term 360.

-358x=(1)/(8)*360^(45)

Reduce the expression by removing the common factor of 8 in the denominator of the first term (1)/(8) and the second term 360.

-358x=1*45

Multiply 1 by 45 to get 45.

-358x=45

Divide each term in the equation by -358.

-(358x)/(-358)=(45)/(-358)

Move the minus sign from the denominator to the front of the expression

-(-(358x)/(358))=(45)/(-358)

Cancel the common factor of 358 in -(358x)/(358).

-(-(358x)/(358))=(45)/(-358)

Remove the common factors that were cancelled out.

-(-x)=(45)/(-358)

Multiply - by the -x inside the parentheses.

-*-x=(45)/(-358)

Multiply - by -x to get x.

x=(45)/(-358)

Move the minus sign from the denominator to the front of the expression.

x=-((45)/(358))

Multiply -1 by the (45)/(358) inside the parentheses.

Answer is:
x=-(45)/(358)